(10%) Problem 6: Two power lines, line 1 and line 2, both of length LL-58 m, are strung east-west between two towers. line lCan you please help with parts a - d
Answer #1

Part A.

B21 = Magnetic field produced by line 1 at line 2

SInce given that r12 <<< LL, So We can treat wires as infinite length wire

Now Magnetic field due to infinite length wire at a distance r12 from wire is given by:

B21 = u0*i1/(2*pi*12)

i1 = 81 A towards west

r12 = 2.5 m

So,

B21 = 4*pi*10^-7*81/(2*pi*2.5)

B21 = 6.48*10^-6 T

Part B.

From Right hand Rule, direction is towards south

Part C.

Now Magnetic force between two current carrying wires is given by:

F21 = i2*L*B21

i2 = 81 A

L = 58 m

B21 = 6.48*10^-6

So,

F21 = 81*58*6.48*10^-6

F21 = 0.03044 N = 30.44*10^-3 N

Part D.

Now We need current in both lines when magnetic force is balanced by force of gravity

Fm = Fg

i*L*B21 = m*g

B21 = u0*i/(2*pi*r12)

M = mass of wire = 890 kg/1000 m = 0.890 kg/m

mass of 58 m wire will be = M*L = (0.890 kg/m)*58 m = 51.62 kg

g = 9.81 m/s^2

So,

i*L*u0*i/(2*pi*r12) = m*g

i^2 = 2*pi*r12*m*g/(u0*L)

Using known values:

i = sqrt (2*pi*2.5*51.62*9.81/(4*pi*10^-7*58))

i = 10446.83 Amp

Since we need how many times larger than initial given current, So

i/iL = 10446.83/81

i/iL = 128.97 Amp = 129 times larger

Please Upvote.